# ComFree-Sim: A GPU-Parallelized Analytical Contact Physics Engine for Scalable Contact-Rich Robotics Simulation and Control

### 12 Mar 2026 v1

@@ I INTRODUCTION
A rigid-body simulation step typically includes collision detection, contact resolution, and time integration [4, 16]. Among these, contact resolution, which computes contact impulses and post-contact velocities, most strongly impacts physical fidelity, numerical stability, and runtime [11]. Main approaches for contact resolution enforce non-penetration and Coulomb friction via the complementarity constraints or equivalent constrained optimization [31, 3, 34], resulting in per-step iterative solves whose cost often grows superlinearly with contact counts. This scaling is prohibitive for batched, differentiable simulation, and online planning. Even state-of-the-art GPU-parallelized phsyics engines, e.g., MJWarp [10] can exhibit substantial step-time growth in dense contact scenes, complicating real-time deployment.
@@
  • Unified 6D analytical contact modeling. We extend complementarity-free point contact model to a 6D formulation that captures tangential, torsional, and rolling friction within an analytical dual-cone resolution framework.
  • GPU parallelization with stable impedance heuristics. We develop GPU implementation for the analytical contact resolution and introduce an effective dual-cone impedance heuristic. This preserves a lightweight user-facing parameterization while also supporting learning-based dynamic impedance scheduling [37].
@@
Together, these results suggest that analytical, decoupled contact resolution can provide a highly competitive and scalable alternative to iterative complementarity-based backends, particularly for the high-frequency, contact-rich rollouts that underpin modern learning and control.
@@ III ComFree-Sim: Analytical Contact Physics Computation and GPU Implementation
Complementarity-free (ComFree) contact modeling was introduced in [14] and subsequently demonstrated in contact-rich control and model learning [40, 37, 41]. Our ComFree-Sim will extend it with (1) a unified 6D contact model covering tangential, torsional, and rolling friction, (2) effective dual-cone impedance heuristics, and (3) GPU parallel implementation [24] as a drop-in backend alternative to MJWarp [10].
@@ ACKNOWLEDGMENT
We thank the MuJoCo Warp (MJWarp) team at Google DeepMind and NVIDIA for making the code publicly available. We also thank Vamsi Sai Abhijit Tadepalli of the IRIS Lab for maintaining the vision-tracking module used in our real-world in-hand manipulation experiments.

### 14 Mar 2026 v2

@@ I INTRODUCTION
A rigid-body simulation step typically includes collision detection, contact resolution, and time integration [4, 16]. Among these, contact resolution, which computes contact impulses and post-contact velocities, most strongly impacts physical fidelity, numerical stability, and runtime [11]. Main approaches for contact resolution enforce non-penetration and Coulomb friction via the complementarity constraints or equivalent constrained optimization [31, 3, 34], resulting in iterative per-step solutions whose cost often grows superlinearly with contact counts. This scaling is prohibitive for batched, differentiable simulation and online planning. Even state-of-the-art GPU-parallelized phsyics engines, e.g., MJWarp [10] can exhibit substantial step-time growth in dense contact scenes, complicating real-time deployment.
@@
  • Unified analytical 6D frictional contact modeling. We extend the complementarity-free point contact model to a 6D formulation that captures tangential, torsional, and rolling friction within an analytical dual-cone contact resolution framework.
  • GPU parallelization with stable impedance heuristics. We develop GPU implementation for the analytical contact resolution and introduce an effective dual-cone impedance heuristic. This preserves a lightweight user-facing parameterization while also supporting learning-based dynamic impedance adaptation [37].
@@
Together, these results suggest that analytical, lightweight contact resolution can provide a highly competitive and scalable alternative to iterative complementarity-based backends, particularly for the high-frequency, contact-rich rollouts that underpin modern learning and control.
@@ III ComFree-Sim: Analytical Contact Physics Computation and GPU Implementation
Complementarity-free (ComFree) contact modeling was proposed in [14] and subsequently demonstrated in contact-rich control and learning [40, 37, 41]. Our ComFree-Sim will extend it with (1) a unified 6D contact model covering tangential, torsional, and rolling friction, (2) effective dual-cone impedance heuristics, and (3) GPU parallel implementation [24] as a drop-in backend alternative to MJWarp [10].
@@ ACKNOWLEDGMENT
We thank the MuJoCo Warp (MJWarp) team at Google DeepMind and NVIDIA for making the code publicly available. We also thank Vamsi Sai Abhijit Tadepalli from the IRIS Lab for maintaining the vision-tracking module used in our real-world in-hand manipulation experiments.